• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M Friend

    Log entry time 02:15:11 on October11,2009

    Entry number 294337

    keyword=Comton 15uA run summary

    Requested 15uA by closing the slits

    Moved electron detector to -2.0mm.

    Run 20301: Electron only run, prescales 100%, FADC trig on edet, 3 plane coincidences. Beam halo is at 12, so we will move to 1.0mm and take coincidence run.

    Edet moved to 1.0mm

    Run 20302: 15uA coincidence run, prescales all at 100%, FADC trig back to the PMT, still 3 plane coincidences. Background is 500Hz/uA, signal is 1200Hz/uA at HV = -1800. Run ended automatically after 20 minutes because of the size limit.

    Edet moved back to garage.